OpenStudy (anonymous):

∫(3x-2)/(x^2(x-1))dx|(2,∞)

OpenStudy (anonymous):

do you know partial derivative?

OpenStudy (bahrom7893):

no this is just improper integration

OpenStudy (anonymous):

then the ans to your qn is ln2

OpenStudy (anonymous):

i did but you have to know the partial derivative i will take steps by steps

OpenStudy (anonymous):

lets deal with number first forget the integral sign

OpenStudy (anonymous):

since this cant be done direct here is set up a/x+bx/x^2+c/[x-1] =[3x-2]/[x^2*(x-1)] then find the value of a b and c

OpenStudy (anonymous):

the values are a=2 b=-3 c=0 you can solve simultaneous eqn or matrix

OpenStudy (anonymous):

now we can solve integration which look like this 2/x-3x/x^2+0/[x-1] this very simple to you now right?

OpenStudy (anonymous):

2lnx-3lnx boundry [2,infinity]

OpenStudy (anonymous):

0-[2ln2-3ln2]=-ln2/3ln2 property if logarithm this is = -ln2^2-3=-ln2^-1=ln2 final ans

OpenStudy (anonymous):

that is how you do, any question facebook me if am not online

OpenStudy (anonymous):

ok thanks a lot!!!